Determine the electric flux for Gaussian surface that contains 100 trillion electrons.

1
Expert's answer
2022-03-14T09:45:38-0400

Answer

Flux is given

"\phi=\frac{Q}{\epsilon_0}\\=\frac{10^{12}*1.6*10^{-19}}{8.85*10^{-12}}\\=180.79*10^2Wb"